OpenStudy (anonymous):

Solve x2 + 9x + 9 = 0

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

we all do

OpenStudy (anonymous):

how so?

OpenStudy (amistre64):

i find complete the square to be the most non traumatic way

OpenStudy (anonymous):

not if middle term is odd is it? you can use formula as well

OpenStudy (anonymous):

[[x = -9/2+(3/2)*sqrt(5)], [x = -9/2-(3/2)*sqrt(5)]]

OpenStudy (anonymous):

although i do like completing the square. i will let amistre write it that way, i will use the formula \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] with \[a=1,b=9,c=9\] to get \[x=\frac{-9\pm\sqrt{9^2-4\times 9\times 1}}{2}\]or \[x=\frac{-9\pm\sqrt{45}}{2}\] \[x=\frac{-9\pm3\sqrt{5}}{2}\]

OpenStudy (amistre64):

x2 + 9x +9 = 0 ; -9 -9 -9 --------------- x^2 +9x + ___ = -9+___ ; complete the square; (b/2)^2 (9/2)^2 = 81/4 x^2 +9x + 81/4 = -9(4)/(4)+81/4 ; compact (x+9/2)^2 = (81-36)/4 = 45/4 we need +-sqrt(45/4) so subtract out the 9/2 \[x = -\frac{9}{2}\pm\frac{\sqrt{45}}{2}\]

OpenStudy (anonymous):

there two methods all worked out!

OpenStudy (anonymous):

awesome!

OpenStudy (anonymous):

try this one if ud like If a baseball player hits a baseball from 4 feet off the ground with an initial velocity of 64 feet per second, how long will it take the baseball to hit the ground? Use the equation h = -16t^2 + 64t + 4

OpenStudy (amistre64):

its a fly ball :)

OpenStudy (anonymous):

lol

OpenStudy (amistre64):

divide out the common of -4 to get: 4t^2 -16t -1

OpenStudy (anonymous):

i got 2+4\[\sqrt{17}\]

OpenStudy (amistre64):

and lets work it :) 4(t^2-4t+4) = 1 + 4(4) (t-2)^2 = 17 t = 2 +- sqrt(17)/2 if i see it right

OpenStudy (amistre64):

t = 2 + sqrt(17)/2 should do it http://www.wolframalpha.com/input/?i=4t^2+-16t+-1